6 years ago Bob was 4 times as old as Peri. In 9 years he will be 3/2 as old.
Answer by stanbon(75887)

6 years ago Bob was 4 times as old as Peri. In 9 years he will be 3/2 as old.
----
Equation:
(b-6) = 4(p-6)
(b+9) = (3/2)(p+9)
========================
Rearrange:
b-4p = -24+6
2b+18 = 3p+27
---
Modify:
b - 4p = -18
2b -3p = 9
---
Finally:
2b - 8p = -36
2b - 3p = 9
----
Subtract top from bottom and solve for "p":
5p = 45
p = 9 (Peri's age now)
---
Solve for "b":
b - 4p = -18
b -4*9 = -18
b = -18+36
b = 18 (Bob's age now)
=========================
Cheers,
Stan H.
==========